OpenStudy (nicoledonley):

Can someone please help me factor the polynomial? 5x^2+8x-4

OpenStudy (triciaal):

to factor set = 0

OpenStudy (nicoledonley):

so would it be 5x^2+8x-4=0 and then it would 5x^2+8x=4 ??

OpenStudy (ac3):

(5x-2)(x+2)

OpenStudy (ac3):

final answer

OpenStudy (triciaal):

standard format ax^2 +bx + c = 0 what are the factors of (a*c) that add up to b? pay attention to the signs. split "b" into its components and rewrite the original. now have 2 groups of 2 terms and take out what is common in each group. each group will have a common factor

OpenStudy (ac3):

wait i think that's wrong

OpenStudy (triciaal):

@Nicoledonley identify your a,b, and c

OpenStudy (ac3):

no i'm right it checks out

OpenStudy (nicoledonley):

a=5x^2 b=8x c=-4 ?

OpenStudy (triciaal):